OpenStudy (anonymous):

What is the solution of the equation? 1+sqrt 3x+3 = sqrt 7x+2

OpenStudy (anonymous):

\[1+ \sqrt{3x+3}=\sqrt{7x+2}\]

OpenStudy (aravindg):

take square root 3x+3 to RHS and square both sides lets see what you get